There are 3 identical boxes containing respectively 1 white and 3 red balls; 2 white and 1 red balls; 4 white and 3 red balls. One box is chosen at random and 2 balls are drawn:
(i) Find the probability that the balls are white and red ?
(ii) If the balls are white and red, what is the probability that they are from the second box?

Grade:12th pass

1 Answers

wdfghyh
5 years ago
The probility that the balls are white and red :
           1 if it is 1st box  then:
                  white=1/13
                   red=3/13
          2 if it is 2nd box then:
                   white=2/13
                   red=1/13
           3 if it is 3rd box then:
                white=4/13
                red=3/13
ANS-2. if they are taken from second box:
                      white=2/13
                       red= 1/13